OpenStudy (anonymous):

Teresa is organizing her closet. She wants to place all of her fifteen pairs of shoes on shelves. If she wants to have the same number of pairs on each shelf, how many choices for the number of shelves does she have? 1

OpenStudy (ganpat):

yup.. u r right, i guess..

OpenStudy (anonymous):

is this multiple choice?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

ok wat r the choices?

OpenStudy (anonymous):

1, 4,5 ,15

OpenStudy (anonymous):

5

OpenStudy (anonymous):

OKAY

OpenStudy (ganpat):

@Cemantha : 5 ?? can u explain ??

OpenStudy (anonymous):

i say 5 because she wants the same number of shoes on each shelve and if she has 5 shelves she will have 3 pairs of shoes on each shelve

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

i thnk i read it wrong

OpenStudy (anonymous):

let me get someone else to come n help...

OpenStudy (anonymous):

@dpaInc can you help

OpenStudy (ganpat):

@Cemantha : it looks little logical.. you can try dividing 15..

OpenStudy (ganpat):

15/1 = 15 15/3 = 5 15/5 = 3 15/15 = 1.. so she has 4 possible options...

OpenStudy (anonymous):

ok then its 4

OpenStudy (ganpat):

answer is 4.. affirmative..

OpenStudy (anonymous):

OKAY